I am having a hard time with this one can someone please help me?
Its a multiple choice question but I am having a hard time.
Solve by using the quadratic formula. 3x^2 = 11x + 4
A) -1/3,4
B) 1/3,-4
C) -1/4,3
D) 1/4,-4

Answer by stanbon(75887)

Solve by using the quadratic formula.
3x^2 - 11x - 4 = 0
x = [11 +- sqrt(121-4*3*-4)]/6
x = [11 +- sqrt(169)]/6
x = [11 +- 13]/6
x = 24/6 = 4 or x = -2/6 = -1/3
=========================
Cheers,
Stan H.